PLEASE SOME ONE HELP ME IF CAN PLEASE
IF(K+1/K)^2=16,THEN K^2+1/K^2=

Answer by ewatrrr(24785)

(K+1/K)^2=16,
K^2+1/K^2= ± 4 considering -4 is an Extraneous Root (K^2 cannot be minus)
K^2 + 1 = 4K^2
1/3 = K^2
√3/3 = K